OpenStudy (anonymous):

A person hired a firm to build a CB radio tower. The firm charges $100 for labor for the first 10 feet. After that, the cost of the labor for each succeeding 10 feet is $25 more than the preceding 10 feet. That is, the next 10 feet will cost $125, the next 10 feet will cost $150, etc. How much will it cost to build a 90-foot tower?needs to be written out how the answer was gotten

OpenStudy (anonymous):

10x100 + 10x125 + 10x150 + 10x175 + 10x200 + 10x225 + 10x250 + 10x275 +10x300 = 10(100 + 125+150+175+200+225+250+300) = 10(1525) = 15250 Hope this answers yr question.......

OpenStudy (anonymous):

if satisfied, pls click the GOOD ANSWER button to show your appreciation and to mark the question as solved and closed :))